# FastGS: Training 3D Gaussian Splatting in 100 Seconds

### v1

@@ 1 Introduction
Novel view synthesis (NVS) is a fundamental problem in computer vision and graphics, with broad applications in augmented reality [zhou2018stereo], virtual reality [xu2023vr] and autonomous driving [wu2023mars]. Neural Radiance Field (NeRF) [mildenhall2021nerf] methods model scenes as continuous volumetric functions and render photorealistic views, but require hours of training per scene. Recently, 3D Gaussian Splatting (3DGS) [kerbl3Dgaussians] has achieved rendering quality comparable to NeRF while offering significantly faster training and rendering speed. It models 3D scenes via explicit Gaussian primitives and employs a tile-based rasterizer. Benefiting from its efficiency, 3DGS has been successfully applied to a wide range of tasks, including dynamic scene reconstruction, surface reconstruction and simultaneous localization and mapping (SLAM). However, a major bottleneck in its current practical application is the extended training time, which often requires tens of minutes per scene, hindering user-friendly deployment.
@@
A detailed analysis of the vanilla 3DGS training pipeline reveals two primary limitations: (1) its adaptive density control (ADC) of Gaussians often introduces a large number of redundant Gaussians, and (2) inefficiencies in the vanilla 3DGS rendering pipeline. While recent works [mallick2024taming, hanson2025speedy, feng2025flashgs, gui2024balanced] have significantly optimized the rendering pipeline, ADC remains a major area for improvement.
@@
ADC in vanilla 3DGS comprises two main components. The first is Gaussian densification, which clones or splits a Gaussian based on its positional gradient. The second is Gaussian pruning, which removes Gaussians with low opacity or oversized scales. Existing 3DGS acceleration methods [fang2024mini, girish2024eagles, papantonakis2024reducing, lan20253dgs2, wang2024adr, deng2024efficient, mallick2024taming, hanson2025speedy, chen2025dashgaussian] have introduced improvements to ADC. One direction of improvement focuses on designing mechanisms to constrain Gaussian densification, aiming to minimize the growth of redundant Gaussians. For example, Taming-3DGS [mallick2024taming] employs a budget-constrained optimization to control Gaussian growth. Similarly, DashGaussian [chen2025dashgaussian] leverages an adaptive Gaussian primitive budgeting method to maintain continuous densification throughout training. The other direction involves refining the pruning strategy to accelerate training by deleting a greater number of redundant Gaussians. For example, Speedy-Splat [hanson2025speedy] applies a soft pruning strategy during densification and a hard pruning strategy afterward.
@@
One major drawback of these methods is the limited effectiveness of their densification and pruning strategies, which fail to maintain rendering quality while avoiding excessive Gaussian redundancy, resulting in an inefficient representation, as illustrated in Fig. 2. This indicates that their Gaussian control strategies are suboptimal. Based on our observations, some densification and pruning methods [kim2024color, kerbl3Dgaussians, kheradmand20243d, ye2024absgs, chen2025dashgaussian, ali2024trimming, girish2024eagles, salman2024elmgs] do not leverage multi-view consistency, while others [zhang2024pixel, mallick2024taming, hanson2025speedy, hanson2025pup] exploit it suboptimally. Specifically, they enforce multi-view consistency merely through Gaussian-related scores, which we argue is insufficient. It may lead to the excessive growth of redundant Gaussians that provide only marginal improvements to the rendering quality from a few viewpoints while contributing little to others. On one hand, for certain densification methods such as Taming-3DGS [mallick2024taming], Gaussian importance is considered across views. However, it fully relies on Gaussian-related scores rather than their actual contribution to rendering quality, resulting in weak multi-view constraints and leading to redundancy. Moreover, it lacks a dedicated redesign of the pruning strategy. On the other hand, for some pruning methods such as Speedy-Splat [hanson2025speedy], multi-view information is also considered, but it uses the gradients of Gaussian rather than evaluating each Gaussian’s contribution to multi-view rendering quality. This indirect enforcement of multi-view consistency leads to significant degradation in rendering quality.
@@
To address the above issues, we propose FastGS, a new, simple, and general 3DGS acceleration framework, capable of training a scene within 100 seconds while maintaining comparable rendering quality, as shown in Fig. 1. In fact, nearly every Gaussian primitive participates in rendering the same region across multiple viewpoints. Our insight is similar to the concept behind bundle adjustment in traditional 3D reconstruction, where each 3D Gaussian should maintain multi-view consistency. This implies that the 3D Gaussian should enhance rendering quality across multiple views of the same region. Therefore, we introduce a multi-view consistent densification (VCD) strategy, which uses a multi-view reconstruction quality importance score to evaluate whether a Gaussian contributes beneficially to the improvement of multi-view rendering quality. Based on the same idea, we propose a multi-view consistent pruning (VCP) strategy, which removes redundant Gaussians that are useless to multi-view rendering quality. Notably, because VCD and VCP accurately identifies which Gaussians need to be densified or pruned, our method does not require a budget mechanism, making it applicable to other tasks easily. To summarize our contributions,
@@
    New, simple and general framework for 3DGS acceleration that can train a scene within 100 seconds while achieving the comparable rendering quality.
@@
    Efficient densification and pruning strategy strictly controls the addition and removal of each Gaussian based on its contribution to multi-view reconstruction quality, greatly accelerating the training process.
@@
    State-of-the-Art Performance on extensive experiments across Mip-NeRF 360, Tanks & Temples, and Deep Blending datasets. Our method outperforms the state-of-the-art (SOTA) approaches in training efficiency while maintaining comparable rendering quality.
  • 4.
    Strong generality across various tasks, such as dynamic scene reconstruction, surface reconstruction, sparse-view reconstruction, large-scale reconstruction and SLAM. FastGS achieves an average 2–7× training speedup.
Figure 2: Gaussian Count Over Training Iterations. Benefiting from the efficient VCD and VCP strategies, FastGS maintains the smallest number of Gaussians throughout the entire training.
@@
2 Related Work
Gaussian densification.
A key bottleneck of 3DGS is the excessive Gaussians that significantly slow down the training process. Recent works [kim2024color, rota2024revising, mallick2024taming, ye2024absgs, zhang2024pixel, chen2025dashgaussian, kheradmand20243d] attempt to address this issue by refining the densification strategy to control the primitive count. Specifically, 3DGS-MCMC [kheradmand20243d] adopts a Monte Carlo strategy for Gaussian densification. Taming-3DGS [mallick2024taming] employs a budgeting mechanism to control Gaussian growth, primarily computing Gaussian scores based on Gaussian-associated properties. DashGaussian [chen2025dashgaussian] also uses a budgeting method to guide Gaussian densification. Nevertheless, they still require millions of Gaussians to maintain rendering quality, resulting in heavy optimization overhead.
Gaussian Pruning.
Besides modifying densification, some methods [ali2024trimming, salman2024elmgs, hanson2025speedy, hanson2025pup, fan2024lightgaussian, girish2024eagles, fang2024mini, papantonakis2024reducing] achieve acceleration by designing pruning strategies to remove a large number of Gaussians. Some of them design importance scores according to Gaussian properties to guide pruning, discarding less critical Gaussians [ali2024trimming, fan2024lightgaussian, girish2024eagles]. Mini-Splatting [fang2024mini] removes a large number of Gaussians through a simplification strategy based on intersection preserving and sampling. PUP 3DGS [hanson2025pup] and Speedy-Splat [hanson2025speedy] remove Gaussians by computing a Gaussian-associated Hessian approximation across all training views. Nonetheless, some of them fail to fully remove redundant Gaussians, while others remove many Gaussians at the cost of significantly degraded rendering quality.
Other Methods.
Some works focus on optimizing 3DGS rasterization or optimization strategies. Taming-3DGS [mallick2024taming] replaces per-pixel with per-splat parallel backpropagation, which significantly speeds up the optimization process and serves as a strong baseline for subsequent research. StopThePop [radl2024stopthepop], FlashGS [feng2025flashgs], and Speedy-sSplat [hanson2025speedy] use precise tile intersection to reduce Gaussian-tile pairs and accelerate rasterization. 3DGS-LM [hollein20243dgs] replaces Adam with Levenberg-Marquardt for faster convergence, and 3DGS² [lan20253dgs2] achieves near second-order convergence via prioritized per-kernel updates.
3 Background: 3D Gaussian Splatting
To construct this representation, a sparse point cloud obtained from SfM is used to initialize the positions of Gaussian primitives. Each primitive $\mathcal{G}_{i}$ is parameterized by mean $\boldsymbol{\mu}_{i}\in\mathbb{R}^{3}$, rotation $r_{i}\in\mathbb{R}^{4}$, scale $s_{i}\in\mathbb{R}^{3}$, opacity $\sigma_{i}\in[0,1)$, and color coefficients $\mathbf{c}_{i}\in\mathbb{R}^{16\times 3}$ represented in view-dependent spherical harmonics (SH). The rotation and scale together define the covariance matrix as
$\mathbf{\Sigma}_{i_{3D}}=R_{i}S_{i}S_{i}^{\top}R_{i}^{\top}.$ (2)
$\boldsymbol{\Sigma}^{\prime}_{i_{3D}}=JW\,\boldsymbol{\Sigma}_{i_{3D}}\,J^{\top}W^{\top},$ (3)
$C(p)=\sum_{i\in N}{c_{i}\alpha_{i}\prod_{j=1}^{i-1}{\left(1-\sigma_{j}\right)}},\quad\alpha_{i}=\sigma_{i}\mathcal{G}_{i}^{\prime}(p),$ (4)
where
@@ 4.1 Overview
The framework of our method is shown in Fig. 3. We initialize 3D Gaussians using SfM point clouds, and train the 3DGS model on multi-view images. The addition and removal of 3D Gaussians are controlled via the proposed multi-view consistent densification and pruning strategies. As illustrated in Fig. 3 and Fig. 3, Taming-3DGS [mallick2024taming] and Speedy-Splat [hanson2025speedy] also estimate Gaussian scores from multiple views, for densification and pruning respectively. However, both rely on Gaussian-associated scores to control the number of Gaussians rather than considering each Gaussian’s contribution to multi-view rendering quality. This suboptimal use of multi-view information results in redundancy for Taming-3DGS [mallick2024taming] and degraded rendering quality for Speedy-Splat [hanson2025speedy] . In contrast, as illustrated in Fig. 3, our method evaluates the importance of each Gaussian based on multi-view reconstruction quality, rather than its properties. Furthermore, our method leverages multi-view consistency for both densification and pruning, which will be detailed in Sections Sec. 4.2 and Sec. 4.3. To further improve rasterization efficiency, Sec. 4.4 introduces the Compact Box (CB) we use.
@@ 4.2 Multi-view Consistent Densification
The vanilla 3DGS [kerbl3Dgaussians] densifies Gaussians solely based on the gradient magnitude in the image space, which leads to a large number of redundant Gaussians. Other densification methods [chen2025dashgaussian, ye2024absgs, fang2024mini] also generate millions of Gaussians, leading to inefficiency. We argue that the redundancy arises because these methods fail to rigorously determine from multiple views whether a Gaussian needs densification. As illustrated in Fig. 3, Taming-3DGS [mallick2024taming] considers multi-view consistency during densification. Howerver, it primarily computes the score based on Gaussian-associated properties (e.g., opacity, scale, depth and gradient) rather than their actual contribution to the rendering quality, making it difficult to enforce strict multi-view consistency for a Gaussian. This also leads to redundancy, as visualized on the left of Fig. 3. Moreover, the computation of its score is complex and relatively inefficient. To address these issues, we propose a new, simple densification strategy VCD based on multi-view consistency. As illustrated in Fig. 3, it computes the average number of high-error pixels in each Gaussian’s 2D footprint across sampled views, where high-error pixels are identified solely from the per-pixel L1 loss between the ground truth and the rendering. As shown in the left of Fig. 3, VCD achieves comparable rendering quality with fewer Gaussians, thereby greatly avoiding redundancy. We then detail how VCD is implemented.
Given $K$ camera views $V=\{v^{j}\}_{j=1}^{K}$, randomly sampled from the training views, together with their corresponding ground-truth images $G=\{g^{j}\}_{j=1}^{K}$ and rendered images $R=\{r^{j}\}_{j=1}^{K}$. For each view $v^{j}$, we compute the error between the rendered color $r^{j,c}_{u,v}$ and the ground-truth color $g^{j,c}_{u,v}$ at pixel $(u,v)$:
$e_{u,v}^{j}=\frac{1}{C}\sum_{c=1}^{C}{\left|r_{u,v}^{j,c}-g_{u,v}^{j,c}\right|},$ (6)
where $c\in\{1,2,\dots,C\}$ denotes the color channel. We then construct the loss map $\mathcal{M}^{j}\in\mathbb{R}^{W\times H}$ from the per-pixel errors:
$\mathcal{M}^{j}=\{\,\hat{e}_{u,v}^{j}\,\}_{u=0,v=0}^{W-1,H-1},$ (7)
where each $\hat{e}_{u,v}^{j}$ is obtained via min–max normalization:
$\hat{e}_{u,v}^{j}=\frac{e_{u,v}^{j}-\min\limits_{u,v}(e_{u,v}^{j})}{\max\limits_{u,v}(e_{u,v}^{j})-\min\limits_{u,v}(e_{u,v}^{j})+\varepsilon},$ (8)
with $\varepsilon$ being a small constant for numerical stability. A threshold $\tau$ is then applied to $\mathcal{M}^{j}$ to identify pixel $p_{h}$ with high reconstruction error, forming a mask:
@@
Next, we need to find the Gaussian primitives associated with these high-error pixels. For each 3D Gaussian primitive $\mathcal{G}_{i}$, we project it onto the 2D image space to obtain its 2D footprint $\Omega_{i}$. We then use an indicator function $\mathbb{I}\big(\mathcal{M}^{j}_{\text{mask}}(p)=1\big)$ to determine whether a pixel has high error. We compute an importance score $s^{+}_{i}$ for each Gaussian primitive, which accumulates the number of high-error pixels contained in the 2D footprint across all sampled views and then averages the accumulated value:
$s_{i}^{+}=\frac{1}{K}\sum_{j=1}^{K}{\sum_{p}^{\varOmega_{i}}{\mathbb{I}\left(\mathcal{M}_{mask}^{j}\left(p\right)=1\right)}},$ (10)
where a higher $s^{+}_{i}$ indicates that the Gaussian consistently lies in high-error regions across multiple views, thus suggesting it as a candidate for densification. A Gaussian primitive $\mathcal{G}_{i}$ is selected for densification if its importance score $s^{+}_{i}$ exceeds a predefined threshold $\tau_{+}$, indicating that it contributes significantly to under-reconstructed regions across multiple views. Notably, we can efficiently determine the number of high-error pixels within the 2D footprint directly from the forward pass of the render.

### v2

@@ 1 Introduction
Novel view synthesis (NVS) is a fundamental problem in computer vision and graphics, with broad applications in augmented reality [zhou2018stereo], virtual reality [xu2023vr], and autonomous driving [wu2023mars]. Neural Radiance Field (NeRF) [mildenhall2021nerf] methods model scenes as continuous volumetric functions and render photorealistic views, but require hours of training per scene. Recently, 3D Gaussian Splatting (3DGS) [kerbl3Dgaussians] has achieved rendering quality comparable to NeRF while offering significantly faster training and rendering speed. It models 3D scenes via explicit Gaussian primitives and employs a tile-based rasterizer. Benefiting from its efficiency, 3DGS has been successfully applied to a wide range of tasks, including dynamic scene reconstruction, surface reconstruction, and simultaneous localization and mapping (SLAM). However, a major bottleneck in its current practical application is the extended training time, which often requires tens of minutes per scene, hindering user-friendly deployment.
@@
A detailed analysis of the vanilla 3DGS [kerbl3Dgaussians] training pipeline reveals two primary limitations: (1) its adaptive density control (ADC) of Gaussians often introduces numerous redundant Gaussians and (2) inefficiencies in the rendering pipeline. While recent works [mallick2024taming, hanson2025speedy, feng2025flashgs, gui2024balanced] have significantly optimized the rendering pipeline, ADC remains a major area for improvement.
@@
ADC in vanilla 3DGS [kerbl3Dgaussians] comprises two main components. The first is Gaussian densification, which clones or splits a Gaussian based on its positional gradient. The second is Gaussian pruning, which removes Gaussians with low opacity or oversized scales. Existing 3DGS acceleration methods [fang2024mini, girish2024eagles, papantonakis2024reducing, wang2024adr, deng2024efficient, mallick2024taming, hanson2025speedy, chen2025dashgaussian, kim2024color, hanson2025pup] have introduced improvements to ADC. One direction of improvement focuses on designing mechanisms to constrain Gaussian densification, aiming to minimize the growth of redundant Gaussians. For example, Taming-3DGS [mallick2024taming] employs a budget-constrained optimization to control Gaussian growth. Similarly, DashGaussian [chen2025dashgaussian] leverages an adaptive Gaussian primitive budgeting method to maintain continuous densification throughout training. The other direction involves refining the pruning strategy to accelerate training by deleting a greater number of Gaussians. For example, Speedy-Splat [hanson2025speedy] applies a soft pruning strategy during densification and a hard pruning strategy afterward.
@@
One major drawback of these methods is the limited effectiveness of their densification and pruning strategies, which fail to maintain rendering quality while avoiding excessive Gaussian redundancy, resulting in an inefficient representation, as illustrated in Fig. 2. This indicates that their Gaussian control strategies are suboptimal. Based on our observations, some densification and pruning methods [kim2024color, chen2025dashgaussian, girish2024eagles, salman2024elmgs] do not leverage multi-view consistency, while others [mallick2024taming, hanson2025speedy, hanson2025pup] exploit it suboptimally. Specifically, they enforce multi-view consistency merely through Gaussian-associated scores, which we argue is insufficient. It may lead to the excessive growth of redundant Gaussians that provide only marginal improvements to the rendering quality from a few viewpoints while contributing little to others. On one hand, for certain densification methods such as Taming-3DGS [mallick2024taming], Gaussian importance is considered across views. However, it fully relies on Gaussian-associated scores rather than their actual contribution to rendering quality, resulting in weak multi-view constraints and leading to redundancy. Moreover, it lacks a dedicated redesign of the pruning strategy. On the other hand, for some pruning methods, such as Speedy-Splat [hanson2025speedy], multi-view information is also considered, but it uses the gradients of Gaussian rather than evaluating each Gaussian’s contribution to multi-view rendering quality. This indirect enforcement of multi-view consistency leads to significant degradation in rendering quality.
@@
To address the above issues, we propose FastGS, a new, simple, and general 3DGS acceleration framework, capable of training a scene in around 100 seconds while maintaining comparable rendering quality, as shown in Fig. 1. In fact, nearly every Gaussian primitive participates in rendering the same region across multiple viewpoints. Our insight is similar to the concept behind bundle adjustment in traditional 3D reconstruction, where each 3D Gaussian should maintain multi-view consistency. This implies that the 3D Gaussian should enhance rendering quality across multiple views of the same region. Therefore, we introduce a multi-view consistent densification (VCD) strategy, which uses a multi-view reconstruction quality importance score to evaluate whether a Gaussian contributes beneficially to the improvement of multi-view rendering quality. Based on the same idea, we propose a multi-view consistent pruning (VCP) strategy, which removes redundant Gaussians that are useless to multi-view rendering quality. Notably, because VCD and VCP accurately identify which Gaussians need to be densified or pruned, our method does not require a budget mechanism, making it easily applicable to other tasks. To summarize our contributions,
Figure 2: Gaussian count over training iterations. Benefiting from the efficient VCD and VCP strategies, FastGS keeps the number of Gaussians consistently low throughout the entire training process on the treehill scene of Mip-NeRF 360 [barron2022mip].
@@
    New, simple, and general framework for 3DGS acceleration that can train a scene in around 100 seconds while achieving comparable rendering quality.
@@
    Efficient densification and pruning strategy strictly controlling the addition and removal of each Gaussian based on its contribution to multi-view reconstruction quality, greatly accelerating the training process.
@@
    General and state-of-the-art performance across various tasks. Our method outperforms state-of-the-art (SOTA) methods in training speed while maintaining comparable rendering quality on static scenes. It generalizes well to dynamic scene reconstruction, surface reconstruction, sparse-view reconstruction, large-scale reconstruction, and SLAM.
2 Related Work
Gaussian Densification. A key bottleneck of 3DGS is the excessive Gaussians that significantly slow down the training process. Recent works [kim2024color, rota2024revising, mallick2024taming, chen2025dashgaussian, deng2024efficient] attempt to address this issue by refining the densification strategy to control the primitive count. Specifically, Taming-3DGS [mallick2024taming] employs a budgeting mechanism to control Gaussian growth, primarily computing importance scores based on Gaussian-associated properties. DashGaussian [chen2025dashgaussian] introduces a resolution-guided primitive scheduler that progressively reconstructs the scene throughout the entire training process. Nevertheless, they still require millions of Gaussians to maintain rendering quality, resulting in heavy optimization overhead.
Gaussian Pruning. Besides modifying densification, some methods [ali2024trimming, salman2024elmgs, hanson2025speedy, hanson2025pup, girish2024eagles, fang2024mini, papantonakis2024reducing] achieve acceleration by designing pruning strategies to remove a large number of Gaussians. Some of them design importance scores according to Gaussian-associated properties to guide pruning, discarding less critical Gaussians [ali2024trimming, girish2024eagles, salman2024elmgs]. Mini-Splatting [fang2024mini] removes a large number of Gaussians through a simplification strategy based on intersection preserving and sampling. PUP 3DGS [hanson2025pup] and Speedy-Splat [hanson2025speedy] remove Gaussians by computing a Gaussian-associated Hessian approximation across all training views. Nonetheless, some of them fail to fully remove redundant Gaussians, while others remove many Gaussians at the cost of significantly degraded rendering quality.
Other Methods. Some works focus on optimizing 3DGS rasterization or optimization strategies. Taming-3DGS [mallick2024taming] replaces per-pixel with per-splat parallel backpropagation, which significantly speeds up the optimization process and serves as a strong baseline for subsequent research. StopThePop [radl2024stopthepop], FlashGS [feng2025flashgs], and Speedy-Splat [hanson2025speedy] use precise tile intersection to reduce Gaussian-tile pairs and accelerate rasterization. 3DGS-LM [hollein20243dgs] replaces Adam with Levenberg-Marquardt for faster convergence, and 3DGS² [lan20253dgs2] achieves near second-order convergence via prioritized per-kernel updates.
3 Background: 3D Gaussian Splatting
To construct this representation, a sparse point cloud obtained from SfM is used to initialize the positions of Gaussian primitives. Each primitive $\mathcal{G}_{i}$ is parameterized by mean $\boldsymbol{\mu}_{i}\in\mathbb{R}^{3}$, rotation $r_{i}\in\mathbb{R}^{4}$, scale $s_{i}\in\mathbb{R}^{3}$, opacity $\sigma_{i}\in\mathbb{R}$, and color coefficients $\mathbf{c}_{i}\in\mathbb{R}^{16\times 3}$ represented in view-dependent spherical harmonics (SH). The rotation and scale together define the covariance matrix as
$\mathbf{\Sigma}_{i_{3D}}=R_{i}S_{i}S_{i}^{\top}R_{i}^{\top}.$ (2)
$\boldsymbol{\Sigma}^{\prime}_{i_{3D}}=JW\,\boldsymbol{\Sigma}_{i_{3D}}\,W^{\top}J^{\top},$ (3)
$C(p)=\sum_{i\in N}{c_{i}\alpha_{i}\prod_{j=1}^{i-1}{\left(1-\alpha_{j}\right)}},\quad\alpha_{i}=\sigma_{i}\mathcal{G}_{i}^{\prime}(p),$ (4)
where
@@ 4.1 Overview
The framework of our method is shown in Fig. 3. We initialize 3D Gaussians using SfM point clouds and train the 3DGS model on multi-view images. The addition and removal of 3D Gaussians are controlled via the proposed multi-view consistent densification and pruning strategies. As illustrated in Fig. 3 and Fig. 3, Taming-3DGS [mallick2024taming] and Speedy-Splat [hanson2025speedy] also estimate importance scores from multiple views, for densification and pruning, respectively. However, both rely on Gaussian-associated scores to control the number of Gaussians. This suboptimal use of multi-view information results in redundancy for Taming-3DGS [mallick2024taming] and degraded rendering quality for Speedy-Splat [hanson2025speedy]. In contrast, as illustrated in Fig. 3, our method evaluates the importance of each Gaussian based on multi-view reconstruction quality, rather than gaussian-associated properties. Furthermore, our method leverages multi-view consistency constraints to effectively guide both densification and pruning, which will be detailed in Sections Sec. 4.2 and Sec. 4.3. To further improve rasterization efficiency, Sec. 4.4 introduces the compact box (CB) we use.
@@ 4.2 Multi-view Consistent Densification
The vanilla 3DGS [kerbl3Dgaussians] densifies Gaussians solely based on the gradient magnitude in the image space, which leads to a large number of redundant Gaussians. Other densification methods [chen2025dashgaussian, ye2024absgs, deng2024efficient] also generate millions of Gaussians, leading to inefficiency. We argue that the redundancy arises because these methods fail to rigorously determine from multiple views whether a Gaussian needs densification. As illustrated in Fig. 3, Taming-3DGS [mallick2024taming] considers multi-view consistency during densification. However, it primarily computes the score based on Gaussian-associated properties (e.g., opacity, scale, depth, and gradient), making it difficult to enforce strict multi-view consistency for a Gaussian. This also leads to redundancy, as visualized on the left of Fig. 3. Moreover, the computation of its score is complex and relatively inefficient. To address these issues, we propose a new, simple densification strategy VCD based on multi-view consistency. As illustrated in Fig. 3, it computes the average number of high-error pixels in each Gaussian’s 2D footprint across sampled views, where high-error pixels are identified solely from the per-pixel L1 loss between the ground truth and the rendering. As shown on the left of Fig. 3, VCD achieves comparable rendering quality with fewer Gaussians, thereby greatly avoiding redundancy. We then detail how VCD is implemented.
Given $K$ camera views $V=\{v^{j}\}_{j=1}^{K}$, randomly sampled from the training views, together with their corresponding ground-truth images $G=\{g^{j}\}_{j=1}^{K}$ and rendered images $R=\{r^{j}\}_{j=1}^{K}$. For each view $v^{j}$, we compute the error between the rendered color $r^{j}_{u,v}$ and the ground-truth color $g^{j}_{u,v}$ at pixel $(u,v)$:
$e_{u,v}^{j}=\frac{1}{C^{\prime}}\sum_{c^{\prime}=1}^{C^{\prime}}{\left|r_{u,v}^{j,c^{\prime}}-g_{u,v}^{j,c^{\prime}}\right|},$ (6)
where $c^{\prime}\in\{1,2,\dots,C^{\prime}\}$ denotes the color channel. We then construct the loss map $\mathcal{M}^{j}\in\mathbb{R}^{W\times H}$ from the per-pixel errors:
$\mathcal{M}^{j}=\mathcal{N}\left(\{\,e_{u,v}^{j}\,\}_{u=0,v=0}^{W-1,H-1}\right),$ (7)
where $\mathcal{N}(\cdot)$ denotes a min–max normalization function. A threshold $\tau$ is then applied to $\mathcal{M}^{j}$ to identify pixel $p_{h}$ with high reconstruction error, forming a mask:
@@
Next, we need to find the Gaussian primitives associated with these high-error pixels. For each 3D Gaussian primitive $\mathcal{G}_{i}$, we project it onto the 2D image space to obtain its 2D footprint $\Omega_{i}$. We then use an indicator function $\mathbb{I}\big(\mathcal{M}^{j}_{\text{mask}}(p)=1\big)$ to determine whether a pixel has high error. We compute an importance score $s^{i}_{d}$ for each Gaussian primitive, which accumulates the number of high-error pixels contained in the 2D footprint across all sampled views and then averages the accumulated value:
$s_{d}^{i}=\frac{1}{K}\sum_{j=1}^{K}{\sum_{p}^{\varOmega_{i}}{\mathbb{I}\left(\mathcal{M}_{mask}^{j}\left(p\right)=1\right)}},$ (9)
where a higher $s^{i}_{d}$ indicates that the Gaussian consistently lies in high-error regions across multiple views, thus suggesting it as a candidate for densification. A Gaussian primitive $\mathcal{G}_{i}$ is selected for densification only when its importance score $s^{i}_{d}$ exceeds a threshold $\tau_{d}$, ensuring that new Gaussians focus on under-reconstructed regions across views. Notably, we can efficiently determine the number of high-error pixels within the 2D footprint directly from the forward pass of the render.